OpenStudy (ineedhelpnow):

what are the steps to solving x/4-1=x/5?

OpenStudy (anonymous):

what are you trying to solve? for x. cause as far as i can see these expression do not equal each other.

OpenStudy (ineedhelpnow):

just says to solve the equation

OpenStudy (akshay_budhkar):

x/4-x/5=1 5x-4x=20 x=20

OpenStudy (anonymous):

this would be how you would solve, but idk, that just doesn't look to equal each other

OpenStudy (akshay_budhkar):

its right.

OpenStudy (anonymous):

as x/4 - x/5 does not = 1 ....everrr

OpenStudy (anonymous):

wait nvm

jimthompson5910 (jim_thompson5910):

Outkast, this is a conditional equation, not an identity.

OpenStudy (anonymous):

it's been a late night

OpenStudy (akshay_budhkar):

then go to sleep!lol

OpenStudy (anonymous):

can't i'm studying for a calculus 2 exam hah

OpenStudy (akshay_budhkar):

you get it ineedhelpnow

OpenStudy (ineedhelpnow):

yes thanks